OpenStudy (karatechopper):

The color with the greatest survival reproduces at a rate of 5 offspring per pair per year, the next highest survival color reproduces at the rate of 4 offspring per year, then 3, 2 and 1 for the remaining phenotypes. Assume that the survival pattern continues indefinitely. Determine how many generations it would take for the population of the most stressed phenotype to drop below 5% of its original portion of the population. The five colors are Red, Turquoise, Dark blue, black, and sage green. For example lets say that sage green was the highest surviving worm phenotype. How would you even

OpenStudy (karatechopper):

Only looking for a process here.

RhondaSommer (rhondasommer):

what process are you even looking for though? like wow thats a lot of words.

OpenStudy (karatechopper):

Trying to see how to calculate when each of the populations would stress out and fall below their original population. So for example lets say that the original population was 203.

OpenStudy (karatechopper):

It's more a mathematical question I think tbh although this is for biology haha. Yes I mean what generation it would take till each phenotype dies off.

OpenStudy (karatechopper):

Dw, I and my peers are extremely confuzzled rn

OpenStudy (baru):

dont we need to know how many pairs of each colour are there to begin with?

OpenStudy (baru):

just to clarify, nothing is dying right? only its percentage share of the population is dropping

OpenStudy (baru):

you could try this: write a differential equation to model population at any time \[dx/dt = k \frac{x}{2}\] x= total population x/2= no of pairs K= growth rate (offsprings per pair per year) solve it \[x_1=C_1e^{k_1 t/2}\]

OpenStudy (baru):

i've added the subscript 1 to indicate that we are looking at one particular colour

OpenStudy (karatechopper):

In my investigation turquoise is the number that is stressed. So we have an original pop of 200 and birds are practically eating the population of worms

OpenStudy (karatechopper):

wait im confused you have two equations?

RhondaSommer (rhondasommer):

poor poor worms

RhondaSommer (rhondasommer):

and things are dying, the birds are eating.

OpenStudy (baru):

i'm sorry, i dont follow the bio part... are there members of the population dying? if so... at what rate?

OpenStudy (karatechopper):

well baru the thing is that our original pop is 200 but for turquoise an average of 57 are picked up by the birds. The birds are the predators so the worms gotta die.

RhondaSommer (rhondasommer):

:( BUT WORMY WAS MY FRIEND...lol

RhondaSommer (rhondasommer):

isnt this what you are doing?

RhondaSommer (rhondasommer):

@karatechopper this is def bio...but i dont remember how to do it

OpenStudy (karatechopper):

Lol that paper you attached is what my lab is

OpenStudy (baru):

hmm.... i dont know any bio. but this looks like a differential equations problem. i dont think you have given all the details in the question. you need to mention what is growing and what is dying at what rate... and how many of each are there to begin with i know only the basics of diff. equations... but if you can make the question more general, i think a lot of people here will help

OpenStudy (karatechopper):

what is d?

OpenStudy (karatechopper):

frequency dying is .42

OpenStudy (karatechopper):

solved by using hardy weinberg

OpenStudy (karatechopper):

so if its dying at .42, the pop would grow at .58 right

OpenStudy (karatechopper):

and we are trying to see when the population will fall below 10. (5% of original populaiton)

RhondaSommer (rhondasommer):

then it wil be an inequality. \[x^y<10\] and x is how many die out to the power of time(y) idk this is a guess...but its an idea...its def an inequality though

RhondaSommer (rhondasommer):

idk 🙅🙅🙅

imqwerty (imqwerty):

so you gotta find out the number of generations after which the population of weakest one falls below 5% of its original portion of the population okaay The five colors are Red, Turquoise, Dark blue, black, and sage green. green is the best and lets say red is the weakest and lets say the population distribution is like this(arranged from weakest to strongest)-> green= \(e\) members black= \(d\) members blue= \(c\) members turquoise = \(b\) members red= \(a\) members % of red in this population= \(\Large \frac{a}{a+b+c+d+e} \times 100\) 5% of this portion is-> \(\Large \frac{5a}{a+b+c+d+e}\) now we will derive some equation for population at \(n^{th}\) generation so if a species "s" has "x" members and each member gives "y" offsprings then total offsprings of 1st generation will be=\(x (y)\) and for second generation it will be-> \((xy)(y)\) so in general we can write that the \(n^{th}\) generation wil have \(x (y)^n\) people note that i have assumed that the older generation die when new generation comes lets say that percentage of red that we want comes in \(n^{th}\) generation so now total population after \(n{th}\) generation will be-\(e(5)^n +d(4)^n+c(3)^n+b(2)^n+a(1)^n\) the percentage of \(red\) is-> \(\Large \frac{a}{d(4)^n+c(3)^n+b(2)^n +e(5)^n} \times 100\) and \(\Large \frac{a}{d(4)^n+c(3)^n+b(2)^n +e(5)^n} \times 100 < \frac{5a}{a+b+c+d+e}\) we can find \(n\) from here but finding n from here might get tedious depending upon the values a,b,..